1. On February 1, 2018, Ellison Co. issued eight-year bonds with a face value of $10,000,000 and a stated interest rate of 9%, payable semiannually on July 1 and January 1. The bonds were sold to yield 10%. The bonds are callable at 101 and convertible.

  1. The issue price of the bonds is
  2. Record the journal entries for February 2018 at issuance and July 1.

2. Using the information above, assume that the bonds issued by Ellison Co. are convertible with each $1,000 convertible into 25 shares of common stock.    Assume that Ellison converts $5,000,000 of bonds on July 1, 2020 into common stock. Prepare the following entries:

a. Entry at February 1, 2018 for issuance of the convertible bonds

b. Entry at July 1, 2020 for the conversion of $5,000,000 of bonds.

3. Using the information above, assume that the remaining bonds are called on December 31, 2020
Answer #1

1.

a.PV of face value of the bonds , (PV @ 5%, 16) = 10,000,000 * 0.45811 = 4581115.2

PV of interest payments (PVAF @5%, 16) = 10,000,000 *4.5% * 10.8377 = 4876996.3

Issue price of the bonds = 4581115.2+ 4876996.3= 9,458,112

b.

Date Account Titles Debit Credit
Feb. 2018 Cash               9,458,112
Unamortized bond discount                  541,888
Bonds Payable 10,000,000
July 1 Interest Expense
[9,458,112 *5%]
                 472,906
Unamortized bond discount                            22,906
Cash
[10,000,000*9%*1/2]
450,000